OpenStudy (anonymous):

cos^6x + sin^6x = 1 - 3 sin^2x + 3 sin^4x LS = RS

OpenStudy (amoodarya):

use (a^3+b^3)=(a+b)(a^2+b^2-ab) a=sin^2x b=cos^2x then simplify

OpenStudy (amoodarya):